# Distributed Model Checking in Graphs Classes

### v1

@@ 2.1 The CONGEST Model
The CONGEST model is a standard model of distributed computation in networks [69]. The network is modeled as a connected simple (i.e., no self-loops nor multiple edges) graphs $G=(V,E)$. The $n\geq 1$ vertices of $G$ model the processing nodes in the networks, and its edges models the communication links. In the following, we may indistinguishably use “node” of “vertex” for referring to the processing elements in the network. Each node is assigned an identifier, which is assumed to be unique in the network. The identifiers may not be between 1 and $n$, but it is assumed that each identifier can be stored on $O(\log n)$ bits in $n$-node networks. We also assume, w.l.o.g., that the nodes know a linear upper-bound on the diameter $D$ of the input graph [69].
Computation proceeds in lockstep, as a sequence of synchronous rounds. All nodes start simultaneously, at the same round. At each round, every node $v$ sends an $O(\log n)$-bit message to each of its neighbors, receives the messages sent by all its neighbors, and performs some individual computation. The round-complexity of an algorithm $A$ is the function $R_{A}(n)=\max_{|V(G)|\leq n}R_{A}(G)$ where $R_{A}(G)$ denotes the number of rounds performed by the nodes executing $A$ in $G$ until all of them terminate.
@@ 2.2 First-Order Formulas on Labeled Graphs
Again, in this paper, all graphs are supposed to be finite, simple (i.e., no self-loops nor multiple edges), and undirected.
@@ 2.2.1 Definition
Recall that a first-order (FO) formula over graphs is a combination (conjunction, disjunction, and negations) of Boolean predicates over variables, which may be quantified or not. Every variable represents a vertex of a graph, and there are two binary predicates: equality $x=y$, and adjacency $\textsf{adj}(x,y)$, where the latter states whether or not there is an edge between the two vertices $x$ and $y$. The variables of an FO formula can be quantified with existential or universal quantifiers. Non-quantified variables are called free variables. Given an FO formula $\varphi$ with free variables $x_{1},\dots,x_{k}$, we make explicit that the latter are free by denoting $\varphi=\varphi(x_{1},\dots,x_{k})$. In the following, FO denotes the set of all first order formulas over undirected graphs.
@@
A graph property is called a first-order property if it can be formulated as the satisfiability of an FO formula without free variables. For instance, existence of a triangle is an FO property, which holds whenever the graph contains a triangle (i.e., a set of three pairwise adjacent nodes). Indeed, this property can be merely formulated as the formula:
$\varphi=\exists x_{1}\exists x_{2}\exists x_{3},\textsf{adj}(x_{1},x_{2})\wedge\textsf{adj}(x_{2},x_{3})\wedge\textsf{adj}(x_{1},x_{3}).$
@@ 2.2.2 FO Formulas for Labeled Graphs
The formalism of FO properties can be extended to labeled graphs. Let $\Lambda$ be a finite set, whose elements are called labels. A $\Lambda$-labeled graph is a pair $(G=(V,E),L)$ where $G$ is an undirected graph, and $L:V\rightarrow 2^{\Lambda}$ is an assignment of a set of labels to each vertex of $G$. An FO formula over $\Lambda$-labeled graphs is a first order formula $\varphi$ over graphs, extended with predicates of the form $\textsf{lab}_{\lambda}(x)$ for every $\lambda\in\Lambda$. The predicate $\textsf{lab}_{\lambda}(x)$ is true if vertex $x$ contains $\lambda$ as one of its labels. Such predicates are called label predicates. Whenever $\textsf{lab}_{\lambda}(x)$ holds, we say that $x$ is labeled with $\lambda$. Note that since $L$ assigns a set of labels to each vertex, it may be the case that $\textsf{lab}_{\lambda}(x)\land\textsf{lab}_{\lambda^{\prime}}(x)$ holds for two different labels $\lambda$ and $\lambda^{\prime}$. It may be convenient to also consider the predicates $\textsf{lab}_{S}$ for all sets $S\subseteq\Lambda$. In this case, $\textsf{lab}_{S}(x)$ holds if and only if $L(x)=S$. In particular, $\textsf{lab}_{S}(x)\land\textsf{lab}_{S^{\prime}}(x)$ is necessarily false whenever $S\neq S^{\prime}$. The set of all first-order logic formulas over $\Lambda$-labeled graphs is denoted $\text{FO}[\Lambda]$.
Given a $\Lambda$-labeled graph $(G,L)$, and a subgraph $G^{\prime}$ of $G$, we abuse notation and denote $(G^{\prime},L)$ the $\Lambda$-labeled graph defined by the restriction of $L$ to the nodes in $G^{\prime}$. For instance, let $\Lambda=\{0,1\}$, and let us consider the following graph property on labeled graphs: “existence of a triangle with a node labeled $1$”. This property can be expressed by the formula:
@@
We may also consider formula $\psi(\overline{x})$ with free variables, such as
@@
The graph $G$ with vertex set $V=\{a,b,c,d\}$ and edge set $E=\{\{a,b\},\{a,c\},\{b,c\},\{c,d\}\}$ with labels $L(a)=\{0\},L(b)=\{0\},L(c)=\{1\}$, and $L(d)=\{1\}$ satisfies $(G,L)\models\varphi$ as the nodes $\{a,b,c\}$ form a triangle where $c$ is labeled $1$. For the same reason, we have that $(G,L)\models\psi(c)$. However, $(G,L)\not\models\psi(a)$ as $a$ is not labeled $1$ (even if it is in a triangle), and $(G,L)\not\models\psi(d)$ as node $d$ is not in a triangle (even if it is labeled $1$).
For $\varphi(x_{1},\dots,x_{k})\in\text{FO}[\Lambda]$, and a $\Lambda$-labeled graph $(G,L)$, we denote by $\textsf{true}(\varphi,G,L)$ the set of all tuples of vertices $(v_{1},\dots,v_{k})$ satisfying that $(G,L)\models\varphi(v_{1},\dots,v_{k})$. In particular, if $\varphi$ has no free variables then $\textsf{true}(\varphi,G,L)$ is equal to the truth value of $(G,L)\models\varphi$. Two $\text{FO}[\Lambda]$ formulas $\varphi_{1}$ and $\varphi_{2}$ are equivalent if $\textsf{true}(\varphi_{1},G,L)=\textsf{true}(\varphi_{2},G,L)$ for every $\Lambda$-labeled graph $(G,L)$. The following basic definitions play a central role in this paper.
  • A formula $\varphi(\overline{x})\in\text{FO}[\Lambda]$ is called quantifier-free if it has no quantifiers, i.e., it can be expressed as a boolean combination of predicates involving only the free variables in $\overline{x}$.
  • A formula $\varphi(\overline{x})\in\text{FO}[\Lambda]$ is existential if it can be written as $\varphi(\overline{x})=\exists\overline{y},\psi(\overline{x},\overline{y})$, where $\psi(\overline{x},\overline{y})$ is quantifier free formula in $\text{FO}[\Lambda]$.
@@
$\varphi(\overline{x})=Q_{1}y_{1}\,\dots\,Q_{t}y_{t}\;\zeta(\overline{x},\overline{y}),$
@@
where $\zeta(\overline{x},\overline{y})\in\text{FO}[\Lambda]$, and $Q_{1},\dots,Q_{t}$ are existential or universal quantifiers. The value of $t$ is the quantifier depth of $\varphi$. Every FO formula admits an equivalent formula in prenex normal form (see, e.g., [20]). This fact can be easily established by induction on the length of the formula, and it can be extended to $\text{FO}[\Lambda]$, i.e., first order formulas for labeled graphs, in a straightforward manner.

### v2

@@ 2 Model and Definitions
In this section, we recall basic concepts of distributed computing, logic, and graphs. Readers familiar with these frameworks may skip the first two sub-sections, but Section 2.3, which introduces and extends the notion of local formulas, and Section 2.4, which recalls properties of graphs with bounded expansion, may require special attention.
@@ 2.1 The CONGEST Model
The CONGEST model is a standard model of distributed computation in networks [57]. The network is modeled as a connected simple (i.e., no self-loops nor multiple edges) graph $G=(V,E)$. The $n\geq 1$ vertices of $G$ model the processing nodes in the networks, and its edges models the communication links. In the following, we may indistinguishably use “node” of “vertex” for referring to the processing elements in the network. Each node is assigned an identifier, which is assumed to be unique in the network. The identifiers may not be between 1 and $n$, but the CONGEST model assumes that each identifier can be stored on $\mathcal{O}(\log n)$ bits in $n$-node networks. In fact, it is usually assumed that all the nodes of a graph $G$ are given a polynomial upper bound $N=\mbox{poly}(n)$ on the number $n$ of vertices in $G$, with all identifiers in $\{1,\dots,N\}$.
Computation proceeds in lockstep, as a sequence of synchronous rounds. All nodes start simultaneously, at the same round. At each round, every node $v$ sends an $\mathcal{O}(\log n)$-bit message to each of its neighbors, receives the messages sent by all its neighbors, and performs some individual computation. The round-complexity of an algorithm $A$ is the function $R_{A}(n)=\max_{|V(G)|\leq n}R_{A}(G)$ where $R_{A}(G)$ denotes the number of rounds performed by the nodes executing $A$ in $G$ until all of them terminate.
As said before, it is worth mentioning the BROADCAST CONGEST model, that is, the weaker variant of the CONGEST model in which, at every round, each node must send the same $\mathcal{O}(\log n)$-bit message to all its neighbors. For the sake of simplifying the notations, we solely refer to the CONGEST model in the statements of our results. Yet, all our upper bound results also holds in the weaker BROADCAST CONGEST model, apart from Theorem 1.4.
Remark.
Throughout the paper, we will use the following convention regarding the inputs and outputs of our algorithms. When an algorithm receives a “global value” as input, whether it be an integer $p$, a logical formula $\varphi$, or a function (e.g., a bound $f:\mathbb{N}\to\mathbb{N}$ on the expansion), we assume that all nodes are given this global value as input. Instead, when an algorithm receives as input a node labeling function $\ell$, we assume that each node $v$ is only given its label $\ell(v)$. Similarly, when an algorithm returns a node-labeling function $\widehat{\ell}$, we mean that each node $v$ returns its label ^$\widehat{\ell}(v)$.
@@ 2.2.1 Definition
Recall that a first-order (FO) formula over graphs is a combination (conjunction, disjunction, and negations) of Boolean predicates over variables, which may be quantified or not. Every variable represents a vertex of a graph, and there are two binary predicates: equality $x=y$, and adjacency $\textsf{adj}(x,y)$, where the latter states whether or not there is an edge between the two vertices $x$ and $y$. The variables of an FO formula can be quantified with existential or universal quantifiers. Non-quantified variables are called free variables. Given an FO formula $\varphi$ with free variables $x_{1},\dots,x_{k}$, we make explicit that the latter are free by denoting $\varphi=\varphi(x_{1},\dots,x_{k})$. In the following, $\mathrm{FO}$ denotes the set of all first order formulas over undirected graphs.
@@
A graph property is called a first-order (FO) property if it can be expressed by a first-order formula with no free variables. For example, the property “the graph contains a triangle” is clearly an FO property: it holds whenever there is a set of three pairwise adjacent vertices in the graph. Indeed, it can be written as the formula
$\varphi\;=\;\exists x_{1}\,\exists x_{2}\,\exists x_{3}\,\bigl{(}\mathrm{adj}(x_{1},x_{2})\;\wedge\;\mathrm{adj}(x_{2},x_{3})\;\wedge\;\mathrm{adj}(x_{1},x_{3})\bigr{)}.$
More generally, the property of containing a fixed graph $H$ as a subgraph (sometimes called “subgraph isomorphism”) is an FO property. Its negation—namely, the property that a graph does not contain $H$ as a subgraph (often called “$H$-freeness”)—is also an FO property, since first-order logic is closed under negation.
@@ 2.2.2 FO Formulas for Labeled Graphs
The formalism of FO properties can be extended to labeled graphs. Let $\Lambda$ be a finite set, whose elements are called labels. A $\Lambda$-labeled graph is a pair $(G,\ell)$ where $G=(V,E)$ is a graph, and $\ell:V\rightarrow 2^{\Lambda}$ is an assignment of a set of labels to each vertex of $G$. An FO formula over $\Lambda$-labeled graphs is a first order formula $\varphi$ over graphs, extended with predicates of the form $\textsf{lab}_{\lambda}(x)$ for every $\lambda\in\Lambda$. The predicate $\textsf{lab}_{\lambda}(x)$ is true if vertex $x$ contains $\lambda$ as one of its labels. Such predicates are called label predicates. Whenever $\textsf{lab}_{\lambda}(x)$ holds, we say that $x$ is labeled with $\lambda$. Note that since $\ell$ assigns a set of labels to each vertex, it may be the case that $\textsf{lab}_{\lambda}(x)\land\textsf{lab}_{\lambda^{\prime}}(x)$ holds for two different labels $\lambda$ and $\lambda^{\prime}$. It may thus be convenient to define the predicates $\textsf{lab}_{S}$ for sets $S\subseteq\Lambda$, as follows. Given $S\in\Lambda$, $\textsf{lab}_{S}(x)$ holds if and only if $\ell(x)=S$. In particular, $\textsf{lab}_{S}(x)\land\textsf{lab}_{S^{\prime}}(x)$ is necessarily false whenever $S\neq S^{\prime}$. The set of all first-order logic formulas over $\Lambda$-labeled graphs is denoted $\mathrm{FO}[\Lambda]$.
Given a $\Lambda$-labeled graph $(G,\ell)$, and a subgraph $G^{\prime}$ of $G$, we abuse notation and denote $(G^{\prime},\ell)$ the $\Lambda$-labeled graph defined by the restriction of $\ell$ to the nodes in $G^{\prime}$. For instance, let $\Lambda=\{0,1\}$, and let us consider the following graph property on labeled graphs: “existence of a triangle with a node labeled $1$”. This property can be expressed by the formula:
@@
We may also consider formula with free variables, such as
@@
The graph $G$ with vertex set $V=\{a,b,c,d\}$ and edge set $E=\{\{a,b\},\{a,c\},\{b,c\},\{c,d\}\}$ with labels $\ell(a)=\{0\},\ell(b)=\{0\},\ell(c)=\{1\}$, and $\ell(d)=\{1\}$ satisfies $(G,\ell)\models\varphi$ as the nodes $\{a,b,c\}$ form a triangle where $c$ is labeled $1$. For the same reason, we have that $(G,\ell)\models\psi(c)$. However, $(G,\ell)\not\models\psi(a)$ as $a$ is not labeled $1$ (even if it is in a triangle), and $(G,\ell)\not\models\psi(d)$ as node $d$ is not in a triangle (even if it is labeled $1$).
For $\varphi(x_{1},\dots,x_{k})\in\mathrm{FO}[\Lambda]$, and a $\Lambda$-labeled graph $(G,\ell)$, we denote by $\textsf{true}(\varphi,G,\ell)$ the set of all tuples of vertices $(v_{1},\dots,v_{k})$ satisfying that $(G,\ell)\models\varphi(v_{1},\dots,v_{k})$. If $\varphi$ has no free variables then $\textsf{true}(\varphi,G,\ell)$ is equal to the truth value of $(G,\ell)\models\varphi$. Two $\mathrm{FO}[\Lambda]$ formulas $\varphi_{1}$ and $\varphi_{2}$ are equivalent if $\textsf{true}(\varphi_{1},G,\ell)=\textsf{true}(\varphi_{2},G,\ell)$ for every $\Lambda$-labeled graph $(G,\ell)$. The following basic definitions play a central role in this paper.
  • A formula $\varphi(\overline{x})\in\mathrm{FO}[\Lambda]$ is called quantifier-free if it has no quantifiers, i.e., it can be expressed as a boolean combination of predicates involving only the free variables in $\overline{x}$.
  • A formula $\varphi(\overline{x})\in\mathrm{FO}[\Lambda]$ is existential if it can be written as $\varphi(\overline{x})=\exists\overline{y}\,\psi(\overline{x},\overline{y})$, where $\psi(\overline{x},\overline{y})$ is quantifier-free formula in $\mathrm{FO}[\Lambda]$.
@@
$\varphi(\overline{x})=Q_{1}y_{1}\,\dots\,Q_{t}y_{t}\;\psi(\overline{x},\overline{y}),$
@@
where $\psi(\overline{x},\overline{y})\in\mathrm{FO}[\Lambda]$, and $Q_{1},\dots,Q_{t}$ are existential or universal quantifiers. The value of $t$ is the quantifier depth of $\varphi$. Every $\mathrm{FO}$ formula admits an equivalent formula in prenex normal form (see, e.g., [21]). This fact can be easily established by induction on the length of the formula, and it can be extended to $\mathrm{FO}[\Lambda]$, i.e., first order formulas for labeled graphs, in a straightforward manner.
Notation.
For every family $\mathcal{G}$ of graphs, and every finite set $\Lambda$, we denote by $\mathcal{G}[\Lambda]$ the set of all $\Lambda$-labeled graphs $(G,\ell)$ with $G=(V,E)\in\mathcal{G}$, and $\ell:V\to 2^{\Lambda}$.
2.3 Local Formulas
Local formulas can be defined as follows. For a graph $G=(V,E)$ and a vertex $v\in V$, let $B_{G}(v,r)$ denote the $r$-neighborhood of $v$ in $G$, that is, the subgraph of $G$ induced by vertices at distance at most $r$ from $v$. By extension, when considering a labeled graph $(G,\ell)$, $B_{G}(v,r)$ is also a labeled graph: each node $u\in B_{G}(v,r)$ is labeled by the same set of labels $\ell(u)$.
Definition 2.1.
Let $\Lambda$ be a finite set of labels, and let $r\geq 0$ be an integer. A formula $\varphi(x)\in FO[\Lambda]$ with one free variable is $r$-local if, for any two $\Lambda$-labeled graphs $(G,\ell)$ and $(G^{\prime},\ell^{\prime})$, and for any two vertices $v\in V(G)$ and $v^{\prime}\in V(G^{\prime})$, we have
$B_{G}(v,r)=B_{G^{\prime}}(v^{\prime},r)\implies\Bigl{(}(G,\ell)\models\varphi(v)\iff(G^{\prime},\ell^{\prime})\models\varphi(v^{\prime})\Bigr{)}.$
where $B_{G}(v,r)=B_{G^{\prime}}(v^{\prime},r)$ means that there exists a label-preserving isomorphism $h:B_{G}(v,r)\to B_{G^{\prime}}(v^{\prime},r)$ such that $h(v)=v^{\prime}$.
We say $\varphi(x)$ is local if there exists $r\geq 0$ such that $\varphi(x)$ is $r$-local.
For technical reasons that will appear clear later, we need to extend the notion of local formulas to formulas $\varphi(x,\overline{y})$ with more than a single free variable. Our notion of locality is defined with respect to the first free variable $x$ of the formula: intuitively, we require that each variable is in a ball centered on $x$. This is different from the notion of Gaifman-locality [29], in which quantified variables are required to be in a ball centered on any of the free variables.
Definition 2.2.
Let $\Lambda$ be a finite set of labels, and let $r\geq 0$ be an integer. A formula $\varphi(x,\overline{y})\in FO[\Lambda]$ is $r$-local if, for any $\Lambda$-labeled graph $(G,\ell)$ with $G=(V,E)$, the following two properties are satisfied:
  • for every pair $(v,\overline{u})\in\textsf{true}(\varphi,G,\ell)$, it holds that $\overline{u}\subseteq B_{G}(v,r)$;
  • for every $\Lambda$-labeled graph $(G^{\prime},\ell^{\prime})$ with $G^{\prime}=(V^{\prime},E^{\prime})$, for every vertices $v\in V,v^{\prime}\in V^{\prime}$, $\overline{u}\subseteq B_{G}(v,r)$ and $\overline{u^{\prime}}\subseteq B_{G^{\prime}}(v^{\prime},r)$, if there exists a label-preserving isomorphism $h:B_{G}(v,r)\to B_{G^{\prime}}(v^{\prime},r)$ such that $h(v)=h(v^{\prime})$ and $h(\overline{u})$ = $h(\overline{u^{\prime}})$, then:
    $(G,\ell)\models\varphi(v,\overline{u})\iff(G^{\prime},\ell^{\prime})\models\varphi(v^{\prime},\overline{u^{\prime}})$
We shall often write $\varphi(\overline{x})$ instead of $\varphi(x,\overline{y})$, in which case it is $x_{1}$ that takes the particular role of $x$ in the definition above.
The above definition is, unfortunately, very difficult in practice, because it provides little information on the form of the considered formula. We therefore define the notion of local form in Definition 2.3. To that end, let us introduce the new predicate $\mathrm{ball}(y,x,r)$ defined as
$G\models\mathrm{ball}(u,v,r)\iff u\in B_{G}(v,r).$
More generally, for $\overline{y}=(y_{1},\dots,y_{k})$, we define
$\mathrm{ball}(\overline{y},x,r)=\bigwedge_{i=1}^{|\overline{y}|}\mathrm{ball}(y_{i},x,r).$
These predicates will be referred to as distance predicates. Note that they can easily be written as existential FO formulas, see Lemma 2.5. We use distance predicates to relativize quantifiers with respect to the $r$-neighborhood of a given vertex (relativization of quantifiers is standard technique, see for example [45, 19]): let us then consider an existential $r$-local formula $\varphi(x)=\exists\overline{y}\,\psi(x,\overline{y})$. Since $\varphi(x)$ is $r$-local, it can be rewritten as
$\varphi(x)=\exists\overline{y}\;\big{(}\psi(x,\overline{y})\wedge\mathrm{ball}(\overline{y},x,r)\big{)}.$
Such a formula, where the locality is enforced explicitly by a distance predicate, is referred to as a formula in $r$-local form. The following definition formalizes this notion.
Definition 2.3.
Let $r\geq 0$ be an integer. A formula $\varphi(x,\overline{y})\in FO[\Lambda]$ with at least one free variable is in $r$-local form if
  • $\varphi(x,\overline{y})=\psi(x,\overline{y})\wedge\mathrm{ball}(\overline{y},x,r)$, where $\psi\in FO[\Lambda]$ is quantifier-free, or
  • $\varphi(x,\overline{y})=\exists z\;\psi(x,\overline{y},z)$, where $\psi\in FO[\Lambda]$ is in $r$-local form, or
  • $\varphi(x,\overline{y})=\mathrm{ball}(\overline{y},x,r)\wedge\forall z\;\big{(}\neg\mathrm{ball}((\overline{y},z),x,r)\vee\psi(x,\overline{y},z)\big{)}$ where $\psi\in FO[\Lambda]$ is in $r$-local form.
We say $\varphi(x,\overline{y})$ is in local form if there exists $r\geq 0$ such that $\varphi(x,\overline{y})$ is in $r$-local form.
When considering a formula with several free variables, we say it is ($r$-)local if it can be rewritten in ($r$-)local form.
Remark.
The last case of Definition 2.3 introduces a lot of redundancy, with two distance predicates, in addition to requiring that $\psi$ must be in $r$-local form. The main motivation for this redundancy is to guarantee that every “subformula” of $\varphi$ remains local. Thanks to this redundancy, all the formulas below are local, despite the fact that negation does not preserve locality for formulas with several free variables. In particular, the following two formulas are equivalent, where the universal quantifier is merely replaced by a negated existential quantifier:
$\mathrm{ball}(\overline{y},x,r)$ $\wedge\forall z\;\big{(}\neg\mathrm{ball}((\overline{y},z),x,r)\vee\psi(x,\overline{y},z)\big{)}$
$\equiv\mathrm{ball}(\overline{y},x,r)\wedge\neg\exists z\;\big{(}\mathrm{ball}((\overline{y},z),x,r)\wedge\neg\psi(x,\overline{y},z)\big{)}.$
Similarly, the formulas
$\exists z\;\big{(}\mathrm{ball}((\overline{y},z),x,r)\wedge\neg\psi(x,\overline{y},z)\big{)},\;\text{and}\;\mathrm{ball}((\overline{y},z),x,r)\wedge\neg\psi(x,\overline{y},z)$
are both local. In other words, the implicit or explicit presence of negations forces us to introduce redundancies in the distance predicates for “balancing” the potential loss of the locality property.
The following result establishes the consistency of Definitions 2.2 and 2.3.
Lemma 2.4.
For every integer $r\geq 0$, every $r$-local formula can be rewritten in $r$-local form.
Proof.
Let $\varphi(x,\overline{y})$ be $r$-local formula, and assume, w.l.o.g, that it is written in prenex normal form $\varphi(x,\overline{y})=Q_{1}z_{1}\,\dots\,Q_{t}z_{t}\;\psi(\overline{x},\overline{y},\overline{z})$. First, we rewrite $\varphi$ by relativizing the quantifier with respect to the $r$-neighborhood of $x$. Intuitively, we replace existential quantification $\exists z$ by $\exists z\in B_{G}(x,r)$, i.e., $z$ must be in the ball of radius $r$ around $x$ in $G$. Similarly universal quantification $\forall z$ is replaced by $\forall z\in B_{G}(x,r)$, i.e., the formula is checked only for nodes $z$ close to $x$. Formally, for the $r$-local formula $\varphi(x,\overline{y})$, we define $\varphi^{r}(x,\overline{y})$ by induction on quantifier depth. If $\varphi$ is quantifier free, then $\varphi^{r}=\varphi$. If $\varphi(x,\overline{y})=\exists z,\zeta(x,\overline{y},z)$, then $\varphi^{r}(x,\overline{y})=\exists z,\mathrm{ball}(z,x,r)\wedge\zeta^{r}(x,\overline{y},z)$. Finally, if $\varphi(x,\overline{y})=\forall z,\zeta(x,\overline{y},z)$, then $\varphi^{r}(x,\overline{y})=\neg\exists z,\mathrm{ball}(z,x,r)\wedge\neg\zeta^{r}(x,\overline{y},z)$.
By definition, for $v\in V$ and $\overline{u}\subseteq B_{G}(v,r)$,
$G\models\varphi^{r}(v,\overline{u})\iff B_{G}(v,r)\models\varphi^{r}(v,\overline{u})\iff B_{G}(v,r)\models\varphi(v,\overline{u}).$
By locality of $\varphi$, we deduce that
$G\models\mathrm{ball}(\overline{u},v,r)\wedge\varphi^{r}(v,\overline{u})\iff B_{G}(v,r)\models\varphi(v,\overline{u})\iff G\models\varphi(v,\overline{u}).$
It only remains to show that the formula $\mathrm{ball}(\overline{y},x,r)\wedge\varphi^{r}(x,\overline{y})$ can be rewritten in $r$-local form. Once again, we proceed by induction on quantifier depth. By definition of $\varphi^{r}$, there are only three cases to consider.
Case 1:
$\varphi$ is quantifier free. Then $\varphi^{r}=\varphi$ and $\mathrm{ball}(\overline{y},x,r)\wedge\varphi(x,\overline{y})$ is already in local form.
Case 2:
$\varphi(x,\overline{y})=\exists z\,\psi(x,\overline{y},z)$. Then,
$\mathrm{ball}(\overline{y},x,r)\wedge\varphi^{r}(x,\overline{y})$ $=\mathrm{ball}(\overline{y},x,r)\wedge\exists z\,\big{(}\mathrm{ball}(z,x,r)\wedge\psi^{r}(x,\overline{y},z)\big{)}$
$=\exists z\,\big{(}\mathrm{ball}((\overline{y},z),x,r)\wedge\psi^{r}(x,\overline{y},z)\big{)}$
By induction, we can rewrite $\mathrm{ball}((\overline{y},z),x,r)\wedge\psi^{r}(x,\overline{y},z)$ in $r$-local form.
Case 3:
$\varphi(x,\overline{y})=\forall z,\psi(x,\overline{y},z)$. Then,
$\mathrm{ball}(\overline{y},x,r)\wedge\varphi^{r}(x,\overline{y})$ $=\mathrm{ball}(\overline{y},x,r)\wedge\neg\exists z\,\big{(}\mathrm{ball}(z,x,r)\wedge\neg\psi^{r}(x,\overline{y},z)\big{)}$
$=\mathrm{ball}(\overline{y},x,r)\wedge\neg\Big{(}\mathrm{ball}(\overline{y},x,r)\wedge\exists z\,\big{(}\mathrm{ball}(z,x,r)\wedge\neg\psi^{r}(x,\overline{y},z)\big{)}\Big{)}$
$=\mathrm{ball}(\overline{y},x,r)\wedge\neg\exists z\,\big{(}\mathrm{ball}((\overline{y},z),x,r)\wedge\neg\psi^{r}(x,\overline{y},z)\big{)}$
$=\mathrm{ball}(\overline{y},x,r)\wedge\forall z\,\big{(}\neg\mathrm{ball}((\overline{y},z),x,r)\vee\psi^{r}(x,\overline{y},z)\big{)}$
$=\mathrm{ball}(\overline{y},x,r)\wedge\forall z\,\Big{(}\neg\mathrm{ball}((\overline{y},z),x,r)\vee\big{(}\mathrm{ball}((\overline{y},z),x,r)\wedge\psi^{r}(x,\overline{y},z)\big{)}\Big{)}$
By induction, $\mathrm{ball}((\overline{y},z),x,r)\wedge\psi^{r}(x,\overline{y},z)$ can be rewritten in $r$-local form.
The study of these three cases completes the proof. ∎
We conclude this section by some simple observation that will be useful at several places in the paper.
Lemma 2.5.
Let $r\in\mathbb{N}$, and let $\overline{x}=(x_{1},\dots,x_{k})$. There exists an existential FO formula
$\beta_{r}(\overline{x})=\exists\overline{y}\,\psi(\overline{x},\overline{y}),$
where $\psi$ is quantifier free such that, for every graph $G$, and every $k$-tuple $\bar{v}$ of vertices of $G$,
$G\models\beta_{r}(\overline{v})\iff G\models\mathrm{ball}(\overline{v},v_{1},r).$
Moreover, for every $r^{\prime}\geq r$, every graph $G$, and every $k$-tuple $\bar{v}$ of vertices of $G$,
$G\models\beta_{r}(\overline{v})\iff G\models\beta_{r,r^{\prime}}(\overline{v}),$
where $\beta_{r,r^{\prime}}(\overline{x})=\exists\overline{y}\,\big{(}\psi(\overline{x},\overline{y})\wedge\mathrm{ball}((\overline{x},\overline{y}),x_{1},r^{\prime})\big{)}$.
Proof.
It is sufficient to describe $\beta_{r}(x,y)$ with two variables only, as $\beta_{r}(x_{1},\dots,x_{k})$ can be written as $\bigwedge_{j=2}^{k}\beta_{r}(x_{1},x_{j})$. We define
$\beta_{r}(x,y)=\exists z_{0}\,\dots\,\exists z_{r}\,\left[(x=z_{0})\wedge\bigvee_{j=0}^{r}\Big{(}(y=z_{j})\wedge\bigwedge_{i=0}^{j-1}\textsf{adj}(z_{i},z_{i+1})\Big{)}\right].$
In this formula, $j$ is the distance between $x$ and $y$, and $z_{0},\dots,z_{j}$ represents the path from $x$ to $y$. This formula satisfies the first item of Lemma 2.5. Let us now prove that it satisfies the second item. Let $r^{\prime}\geq r$. By definition, if $G\models\beta_{r,r^{\prime}}(u,v)$ then $G\models\beta_{r}(u,v)$. Conversely, let us assume that $G\models\beta_{r}(u,v)$. Then, there exists some $\overline{w}=w_{0},\dots,w_{r}$ such that $G\models\psi(u,v,\overline{w})$. Moreover, there exists $j\in[0,r]$ such that
$G\models(u=w_{0})\wedge(v=w_{j})\wedge\bigwedge_{i=0}^{j-1}\textsf{adj}(w_{i},w_{i+1}).$
For every $i\in[0,j]$, let $w^{\prime}_{i}=w_{i}$, and, for every $i\in[j+1,r]$, let $w^{\prime}_{i}=u$. By definition, every vertex $w^{\prime}_{i}$ satisfies $\mathrm{ball}(w^{\prime}_{i},u,r)$, and thus $\mathrm{ball}(w^{\prime}_{i},u,r^{\prime})$ since $r^{\prime}\geq r$. This implies that
$G\models(u=w^{\prime}_{0})\wedge(v=w^{\prime}_{j})\wedge\bigwedge_{i=0}^{j-1}\textsf{adj}(w^{\prime}_{i},w^{\prime}_{i+1})\wedge\mathrm{ball}((u,v,\overline{w^{\prime}}),u,r^{\prime}).$
As a consequence, $G\models\psi(u,v,\overline{w^{\prime}})\wedge\mathrm{ball}((u,v,\overline{w^{\prime}}),u,r^{\prime})$, and thus $G\models\beta_{r,r^{\prime}}(u,v)$. ∎
Lemma 2.6.
Let $\Lambda$ be a set of labels, and let $\varphi(\overline{x})\in\mathrm{FO}[\Lambda]$ be a $r$-local, quantifier-free formula. If $\varphi(\overline{x})=\varphi_{1}(\overline{x})\vee\varphi_{2}(\overline{x})$ for some formulas $\varphi_{1}$ and $\varphi_{2}$, then both $\varphi_{1}$ and $\varphi_{2}$ must be $r$-local.
Proof.
Since $\varphi_{1}$ and $\varphi_{2}$ play symmetric roles, we only prove that $\varphi_{1}$ is local. Since $\varphi$ is quantifier free, and therefore $\varphi_{1}$ as well, checking the $r$-locality of $\varphi_{1}$ is equivalent to checking that, for all $\Lambda$-labeled graphs $(G,\ell)$,
$(G,\ell)\models\varphi_{1}(\overline{v})\implies\overline{v}\subseteq B_{G}(v_{1},r).$
This holds trivially as, by locality of $\varphi$,
$(G,\ell)\models\varphi_{1}(\overline{v})\implies(G,\ell)\models\varphi(\overline{v})\implies\overline{v}\subseteq B_{G}(v_{1},r),$
which completes the proof. ∎